# Rethinking Machine Unlearning for Large Language Models

### v3

@@ 4 LLM Unlearning Methods Revisited: Prior Art and Overlooked Principles
Existing LLM unlearning methods can be broadly categorized into two groups: model-based and input-based. Model-based methods involve modifying the weights and/or architecture components of LLMs to achieve the unlearning objective (Jang et al., 2022; Lu et al., 2022; Yao et al., 2023; Yu et al., 2023; Chen & Yang, 2023; Zhang et al., 2023c; Hase et al., 2023a; Wu et al., 2023b; Rafailov et al., 2023), e.g., following the mathematical formulation in Sec. 3. Input-based methods design input instructions (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023), such as in-context examples or prompts, to guide the original LLM (without parameter updating) towards the unlearning objective.
In the literature, the predominant research emphasis lies on model-based methods as shown in Tab. 1. We review some of the most representative approaches below.
$\bullet$ Gradient ascent and its variants (Jang et al., 2022; Yao et al., 2023; Yu et al., 2023): Gradient ascent (GA) stands as one of the most straightforward unlearning methods, updating the model parameters by maximizing the likelihood of mis-prediction for the samples within the forget set $\mathcal{D}_{\mathrm{f}}$ (Jang et al., 2022). However, it is worth noting that GA alone can be sensitive to the choice of hyperparameters during optimization, such as the number of ascent steps and the learning rate (Jia et al., 2023; Fan et al., 2024). This has given rise to improved variants of GA. For example, Yao et al. (2023) incorporate random labeling to augment the unlearning objective and ensure utility preservation on the retain set $\mathcal{D}_{\mathrm{r}}$.
@@
Another variant of GA is to transform it into a gradient descent approach, which aims to minimize the likelihood of predictions on relabeled forgetting data (Yu et al., 2023). This gradient descent-based fine-tuning, over relabeled or randomly labeled forgetting data, is also employed in (Eldan & Russinovich, 2023), where generic translations are used to replace the unlearned texts. GA and its variants often involve fine-tuning pre-trained LLMs for unlearning purposes. To enhance efficiency, parameter-efficient fine-tuning (PEFT) techniques could be employed. Chen & Yang (2023) fine-tune an adapter over the unlearning objective that acts as an unlearning layer within the LLM. Zhang et al. (2023c) use LoRA to create task vectors and accomplish unlearning by negating tasks under these task vectors.
@@
$\bullet$ Localization-informed unlearning (Meng et al., 2022; Yu et al., 2023; Wu et al., 2023b): The pursuit of parameter efficiency is also in line with the objective of identifying and localizing a subset of model units (e.g., layers, weights, or neurons) that are essential for the unlearning task. For example, the process of localization can be accomplished through representation denoising, also known as causal tracing, in (Meng et al., 2022), focusing on the unit of model layers. Previous work shows that it is important to delete information about unlearning targets wherever it is represented in models in order to protect against attacks (Patil et al., 2024). In addition, gradient-based saliency (Yu et al., 2023) is employed to identify the crucial weights that need to be fine-tuned to achieve the unlearning objective. In (Wu et al., 2023b), neurons that respond to unlearning targets are identified within the feed-forward network and subsequently selected for knowledge unlearning. In the context of vision models, unlearning can also benefit from localizing weights salient to unlearning, as demonstrated by Jia et al. (2023) and Fan et al. (2024). Furthermore, the concept of localization-informed unlearning resonates with the future modular machine learning solution development (Menik & Ramaswamy, 2023). This modularity allows the emerging foundation models to be partitioned into manageable subparts, facilitating easier maintenance and independent updates for each component.
@@
$\bullet$ Influence function-based methods: While the influence function (Koh & Liang, 2017; Bae et al., 2022) is a standard approach to assess the effect of data removal on model performance (Izzo et al., 2021; Warnecke et al., 2021), it is not commonly employed in the context of LLM unlearning for two main reasons: the computational complexity involved in inverting the Hessian matrix, and the reduced accuracy resulting from the use of approximations in influence function derivation (Jia et al., 2023). However, the potential of influence functions in LLM unlearning may be underestimated, given that scalability issue has been improved by Grosse et al. (2023), and approximation errors can be mitigated by focusing on localized weights that are salient to unlearning, as described in the previous category.
@@
$\bullet$ Other model-based methods: Other types of LLM unlearning methods fall outside the categories mentioned above. For example, Jang et al. (2022); Chen & Yang (2023) show that sequential unlearning outperforms batch unlearning However, as indicated by Gu et al. (2024), sequential editing of LLMs may compromise their general capabilities. Therefore, the study of sequential unlearning in LLMs continues to be an unresolved issue.
@@
$\bullet$ Input-based vs. model-based: Input-based strategies (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023) have also shown promise in addressing the challenges posed by the restricted access to black-box LLMs and achieving parameter efficiency of LLM unlearning. Here the learnable parameters are given by input prompts rather than model weights/architecture components. However, we posit that input-based methods may not necessarily yield genuinely unlearned models, leading to weaker unlearning strategies compared to model-based methods because modifying the inputs of LLMs alone may not be sufficient to completely erase the influence of unlearning targets (Toyer et al., 2023). This assertion is also supported by the existence of hard or even adversarial in-scope examples associated with unlearning targets and the challenge to remove their influence in LLMs (Zhong et al., 2023; Patil et al., 2024).
Exploring data-model interactions.
@@
Relationship with model editing.
@@
Despite the aforementioned connections, there are clear distinctions between LLM unlearning and editing. First, the unlearning response is sometimes unknown compared to the editing response. The specificity of an incorrect or improper unlearning response could be seen as a form of LLM hallucination after unlearning. Second, although unlearning and model editing may share some common algorithmic foundations, the former does not create new answer mappings. Rather, its central aim is the comprehensive elimination of the influence attributed to a specific knowledge or concept within a pre-trained LLM. Third, we can differentiate model editing from unlearning from the perspective of ‘working memory’. It is known in (Li et al., 2022) that working memory in LLMs is maintained through neuron activations rather than weight-based long-term memory. Thus, existing memory-based model editing techniques (Li et al., 2022; Mitchell et al., 2022; Madaan et al., 2022; Zheng et al., 2023) focus on updating short-term working memory instead of altering the long-term memory encapsulated in the model’s weights. Yet, we posit that unlearning requires more mechanistic approaches that facilitate ‘deep’ modifications to pre-trained LLMs.
Adversarial training for robust unlearning.
@@
An increasing body of research highlights the weaknesses of existing unlearning methods (Shi et al., 2023; Maini et al., 2024), particularly in their vulnerability to test-time adversaries attempting to jailbreak unlearned models. This issue has been explored by Patil et al. (2024) for LLMs and Zhang et al. (2023d) for diffusion models, and inspires us to integrate adversarial training (Madry et al., 2017) into the unlearning process, resulting in what we term adversarial unlearning. However, this approach has received relatively little attention thus far. While adversarial unlearning increases training costs, it also presents new opportunities. For instance, localization-informed unlearning can significantly reduce the computation expenses associated with adversarial unlearning by focusing on a small portion of model units for updating. Furthermore, advanced adversarial training techniques such as fast adversarial training (Shafahi et al., 2019; Wong et al., 2020; Zhang et al., 2022) and generalized adversarial training (Zhu et al., 2019; Kumari et al., 2019; Zhang et al., 2022; Robey et al., 2023; Casper et al., 2024b) have the potential to enhance the scalability of adversarial unlearning while preserving its effectiveness.
Reinforcement learning.
@@
The mainstream technique for aligning LLMs with human values is RLHF (reinforcement learning from human feedback) and its variants (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Yuan et al., 2023; Lee et al., 2023; Rafailov et al., 2023; Casper et al., 2023). However, RLHF is resource-intense: (1) it requires human inputs that are expensive to collect, and (2) it is computationally costly (i.e., the standard three-stage aligning procedure). LLM unlearning arises as an alternative aligning method, where collecting negative (i.e., low-quality and harmful) samples is much easier through user reporting or (internal) red teaming than positive (i.e., high-quality and helpful) samples which often require hiring humans. Furthermore, we can envision a reinforced unlearning paradigm with a properly defined reward function for the unlearned tasks. In a similar vein, Rafailov et al. (2023); Lu et al. (2022) utilize a reward model to assist in unlearning.
5 Assessment
In this section, we discuss key considerations for designing an effective evaluation pipeline for LLM unlearning.
Tasks and benchmarks.
@@ 5 Assessment
Datasets related to harmful content degeneration, personal identification information removal, and copyrighted information prevention serve as suitable benchmarks for evaluating the effectiveness of LLM unlearning. Some notable examples of these datasets include: The Enron dataset, which comprises employee emails publicly disclosed during Enron’s legal investigation by the Federal Energy Regulatory Commission (Wu et al., 2023b), the Training Data Extraction Challenge dataset used in (Jang et al., 2022), the Harry Potter book series dataset (Eldan & Russinovich, 2023; Shi et al., 2023), the toxicity generation dataset (Lu et al., 2022; Gehman et al., 2020), and the TOFU dataset for fictitious unlearning (Maini et al., 2024).
Evaluation of unlearning effectiveness.
@@
We examine efficacy from three perspectives: comparison with retraining (i.e., the gold standard of unlearning), ‘hard’ in-scope evaluation or robustness, and training data detection.
@@
$\bullet$ LLM unlearning vs. retraining. In classic unlearning paradigms (Golatkar et al., 2020; Thudi et al., 2022; Jia et al., 2023; Fan et al., 2024), retraining a model from scratch after removing the forgotten data from the original training set is regarded as exact unlearning. However, the scalability challenges of retraining LLMs make it difficult to establish a performance upper bound for evaluating LLM unlearning. A recent solution by Maini et al. (2024) is to incorporate fictitious data (synthetic author profiles) into the model training paradigm. Since the injected set never appeared in the original pretraining set, LLM fine-tuning can simulate the retraining process over the newly-introduced set. Despite the progress with regard to the specialized TOFU dataset (Maini et al., 2024), there is still a pressing need to advance current evaluation metrics and pipelines to accurately assess the gap between (approximate) LLM unlearning methods and exact unlearning.
@@
$\bullet$ ‘Hard’ in-scope evaluation or robustness. As demonstrated in Sec. 3, unlearning is generally context and task-dependent, contingent upon an unlearning scope. Another effectiveness metric of LLM unlearning is to ensure forgetting concerning in-scope unlearned examples, even for those ‘hard’ ones that fall within the unlearning scope but may not be directly associated with the unlearning targets. The assessment of ‘hard’ in-scope examples can be achieved by techniques such as paraphrasing what LLMs intend to forget or creating multi-hop questions (Zhong et al., 2023). Evaluating ‘hard’ in-scope examples aligns seamlessly with the underlying principles of ‘worst-case’ or ‘adversarial’ evaluation methods for unlearning, resembling LLM jailbreaking attacks (Zhang et al., 2023d; Yong et al., 2023; Patil et al., 2024). For instance, it is shown by Yong et al. (2023) that unlearning a scope using an English-only example would not guarantee a similar unlearned outcome when translated into other languages. It is also crucial to evaluate the robustness of unlearned LLMs after fine-tuning. Recent studies have revealed that fine-tuning LLMs can sometimes lead to the re-emergence of behaviors that were not anticipated (Yang et al., 2023; Qi et al., 2023; Lermen et al., 2023; Yong et al., 2023).
@@
$\bullet$ Training data detection or membership inference. Membership inference attacks (MIA) (Shokri et al., 2017), designed to detect if a data point is part of a victim model’s training set, serve as a crucial data privacy-unveiled metric for evaluating machine unlearning methods (Thudi et al., 2022; Jia et al., 2023). This metric gains even more significance in the context of LLM unlearning, particularly when retraining is not an option. This concept is also connected to training data memorization (Carlini et al., 2022), as well as training data extraction attacks (Nasr et al., 2023) in LLMs. In the realm of LLM unlearning, privacy-related evaluation metrics have been explored and considered in various studies (Shi et al., 2023; Wu et al., 2023b; Jang et al., 2022; Pawelczyk et al., 2023; Maini et al., 2024).
Utility preservation.
@@
Another crucial metric is to ensure the retained generation capabilities of unlearned LLMs on standard language modeling tasks that fall outside the unlearning scope. For example, evaluation on natural language understanding tasks (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023) and perplexity (Ilharco et al., 2022; Zhang et al., 2023c) has been considered in the literature. In line with evaluating the effectiveness of LLM unlearning on ‘hard’ in-scope examples, it is equally crucial to assess utility preservation using ‘hard’ out-of-scope examples, achieved by e.g., using data transformations. Lastly, we note that it can be difficult to determine the exact scope for some unlearning target (Hase et al., 2023b; Cohen et al., 2023), so part of the challenge here is deciding which generation capabilities should be retained in the first place.
Efficiency.
@@
Computation cost has been a predominant efficiency metric when evaluating LLM unlearning methods, as demonstrated in Table 1. In addition to that, efforts have been made to extend LLM unlearning to black-box models, without access to model parameters, as demonstrated by Pawelczyk et al. (2023). Furthermore, memory efficiency could also serve as a crucial efficiency metric. The distinction from parameter efficiency is that current parameter-efficient fine-tuning methods still impose substantial memory costs for storing LLMs and for executing back-propagation (Malladi et al., 2023). Thus, a future research direction is to explore memory-efficient fine-tuning methods for LLM unlearning.

### v4

@@ 4 Current Unlearning Techniques and Overlooked Principles
Existing LLM unlearning methods can be broadly categorized into two groups: model-based and input-based. Model-based methods involve modifying the weights and/or architecture components of LLMs to achieve the unlearning objective (Jang et al., 2022; Lu et al., 2022; Yao et al., 2023; Yu et al., 2023; Chen & Yang, 2023; Zhang et al., 2023c; Hase et al., 2023a; Wu et al., 2023b; Rafailov et al., 2023), e.g., following the mathematical formulation in Sec. 3. Input-based methods design input instructions (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023; Thaker et al., 2024; Muresanu et al., 2024; Liu et al., 2024), such as in-context examples or prompts, to guide the original LLM (without parameter updating) towards the unlearning objective. In the literature, the predominant research emphasis lies on model-based methods as shown in Table 1. Below, we begin with a review of the most representative approaches for LLM unlearning.
Review of existing unlearning principles
@@
Gradient ascent (GA) and its variants: GA stands as one of the most straightforward unlearning methods, updating the model parameters by maximizing the likelihood of mis-prediction for the samples within the forget set $\mathcal{D}_{\mathrm{f}}$ (Jang et al., 2022; Yao et al., 2023). However, it is worth noting that GA alone can be sensitive to the choice of hyperparameters during optimization (Jia et al., 2023; Fan et al., 2024b), which can lead to unlearning failures such as catastrophic collapse (Zhang et al., 2024b). This has given rise to improved variants of GA. For example, negative preference optimization (NPO) (Zhang et al., 2024b) treats the forgotten data exclusively as negative examples in direct preference optimization (DPO) (Rafailov et al., 2023). This turns the unlearning problem into a minimization problem over the NPO loss, mitigating the issue of catastrophic collapse. Another variant also transforms GA into a gradient descent approach by minimizing the likelihood of predictions on relabeled forgetting data (Yao et al., 2023; Yu et al., 2023). This GA-based fine-tuning, over relabeled forgetting data, is also employed in (Eldan & Russinovich, 2023), where generic translations are used to replace the unlearned texts. GA and its variants often involve fine-tuning pre-trained LLMs for unlearning purposes. To enhance efficiency, parameter-efficient fine-tuning (PEFT) techniques could be employed. For example, an adapter acts as an unlearning layer within the LLM in (Chen & Yang, 2023), and LoRA is used to create task vectors and accomplish unlearning by negating tasks under these task vectors in (Zhang et al., 2023c).
@@
Localization-informed unlearning: The pursuit of parameter efficiency is also in line with the objective of identifying and localizing a subset of model units (e.g., layers, weights, or neurons) that are essential for the unlearning task. For example, the process of localization can be accomplished through representation denoising, also known as causal tracing, in (Meng et al., 2022; Patil et al., 2024), focusing on the unit of model layers. In addition, gradient-based saliency (Yu et al., 2023) is employed to identify the crucial weights that need to be fine-tuned to achieve the unlearning objective. In (Wu et al., 2023b), neurons that respond to unlearning targets are identified within the feed-forward network and subsequently selected for knowledge unlearning. In the context of vision models, unlearning can also benefit from localizing weights salient to unlearning, as demonstrated in (Jia et al., 2023; Fan et al., 2024b). Furthermore, the concept of localization-informed unlearning resonates with the future modular machine learning solution development (Menik & Ramaswamy, 2023). This modularity allows the emerging foundation models to be partitioned into manageable subparts, facilitating easier maintenance and independent updates for each component.
@@
Influence function-based methods: While the influence function (Koh & Liang, 2017; Bae et al., 2022) is a standard approach to assess the effect of data removal on model performance (Izzo et al., 2021; Warnecke et al., 2021), it is not commonly employed in the context of LLM unlearning for two main reasons: the computational complexity involved in inverting the Hessian matrix, and the reduced accuracy resulting from the use of approximations in influence function derivation (Jia et al., 2023). However, the potential of influence functions in LLM unlearning may be underestimated. For example, It has been shown in (Jia et al., 2024) that the influence function approach can be integrated with second-order optimization, transforming static, one-shot influence unlearning into a dynamic, iterative second-order optimization-driven unlearning method, which yields improved unlearning effectiveness. In addition, the approximation error arising from influence function derivation could be mitigated by focusing on localized weights that are salient to unlearning, as described in the previous category.
@@
Outside the categories mentioned above, sequential unlearning (Jang et al., 2022; Chen & Yang, 2023) has been shown to perform better than batch unlearning. However, as indicated by (Gu et al., 2024), sequential editing of LLMs may compromise their general capabilities. Therefore, further studies are needed to better understand and improve sequential unlearning in LLMs.
@@
Input-based vs. model-based: Input-based strategies (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023; Thaker et al., 2024; Muresanu et al., 2024; Liu et al., 2024) have also shown promise in addressing the challenges posed by the restricted access to black-box LLMs and achieving parameter efficiency of LLM unlearning. Here the learnable parameters are given by input prompts rather than model weights/architecture components. However, we posit that input-based methods may not necessarily yield genuinely unlearned models, leading to weaker unlearning strategies compared to model-based methods because modifying the inputs of LLMs alone may not be sufficient to completely erase the influence of unlearning targets (Toyer et al., 2023). This assertion is also supported by the existence of hard or even adversarial in-scope examples associated with unlearning targets and the challenge to remove their influence in LLMs (Zhong et al., 2023; Patil et al., 2024; Lynch et al., 2024). While most input-based methods fall short in this manner, a recently proposed method (Liu et al., 2024), which uses an external classifier as a guardrail and adds learned corruptions to the detected prompts in the embedding space, demonstrates a high resemblance between the resulting unlearned model and the retrained model across multiple tasks and up to 100 models up to over 200B parameters. This suggests that properly crafted inputs could achieve a sense of indistinguishability (between the unlearned and retrained models), albeit in the output space instead of the model space.
Exploring data-model interactions
@@
Relationship with model editing
@@
Despite the aforementioned connections, there are clear distinctions between LLM unlearning and editing. First, the unlearning response is sometimes unknown compared to the editing response. The specificity of an incorrect or improper unlearning response might be seen as a form of LLM hallucination after unlearning. Second, although unlearning and model editing may share some common algorithmic foundations, the former does not create new answer mappings. Rather, its central aim is the comprehensive elimination of the influence attributed to a specific knowledge or concept within a pre-trained LLM. Third, we can differentiate model editing from unlearning from the perspective of ‘working memory’. It is known in (Li et al., 2022) that working memory in LLMs is maintained through neuron activations rather than weight-based long-term memory. Thus, existing memory-based model editing techniques (Li et al., 2022; Mitchell et al., 2022; Madaan et al., 2022; Zheng et al., 2023) focus on updating short-term working memory instead of altering the long-term memory encapsulated in the model’s weights. Yet, we posit that unlearning requires more mechanistic approaches that facilitate ‘deep’ modifications to pre-trained LLMs.
Adversarial training for robust unlearning
@@
An increasing body of research highlights the weaknesses of existing unlearning methods (Shi et al., 2023; Lynch et al., 2024; Patil et al., 2024), particularly in their vulnerability to test-time adversaries attempting to jailbreak unlearned models for unlearned information extraction. This issue has been explored in (Patil et al., 2024; Lynch et al., 2024) for LLMs and in (Zhang et al., 2023d) for diffusion models, and inspires us to integrate adversarial training (Madry et al., 2017) into the unlearning process, resulting in what we term adversarial unlearning. However, this approach has received relatively little attention thus far. To be specific, adversarial unlearning could be formulated as a two-player game (Madry et al., 2017; Zhang et al., 2022), where the defender focuses on LLM unlearning, while the attacker generates jailbreaking attacks aimed at reverse engineering the forgotten information from the model post-unlearning. While adversarial unlearning increases training costs, it also presents new opportunities. For instance, localization-informed unlearning can significantly reduce the computation expenses associated with adversarial unlearning by focusing on a small portion of model units for updating. Furthermore, advanced adversarial training techniques, such as fast adversarial training (Shafahi et al., 2019; Wong et al., 2020; Zhang et al., 2022) and generalized adversarial training in the latent space (Zhu et al., 2019; Kumari et al., 2019; Robey et al., 2023; Casper et al., 2024b), have the potential to enhance the scalability of adversarial unlearning while preserving its effectiveness.
Reinforcement learning and machine unlearning
@@
The mainstream technique for aligning LLMs with human values is RLHF (reinforcement learning from human feedback) and its variants (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Yuan et al., 2023; Lee et al., 2023; Rafailov et al., 2023; Casper et al., 2023). However, RLHF is sometimes resource-intense: (1) it requires human inputs that are expensive to collect, and (2) it is computationally costly (i.e., the standard three-stage aligning procedure). LLM unlearning arises as an alternative aligning method, where collecting negative (i.e., low-quality and harmful) samples is much easier through user reporting or (internal) red teaming than positive (i.e., high-quality and helpful) samples which often require hiring humans. Furthermore, reinforcement learning techniques can be leveraged to assist LLM unlearning, leading to a reinforced unlearning paradigm with a properly defined reward function for the unlearned tasks (Lu et al., 2022). Another example is advancing LLM unlearning using DPO (direct preference optimization) (Rafailov et al., 2023), which simplifies the reinforcement learning part and only requires positive and negative data. The LLM unlearning method NPO (negative preference optimization) (Zhang et al., 2024b) adopts the negative example-only DPO loss as the forget loss, while the PO (preference optimization) method (Maini et al., 2024) introduces targeted unlearning responses such as ‘I don’t know’ or responses stripped of sensitive information, treating these exclusively as positive examples for preference alignment.
5 Assessing LLM Unlearning
@@
There is a pressing need to develop a standardized evaluation pipeline for LLM unlearning. Datasets related to harmful content degeneration, personal identification information removal, and copyrighted information prevention have served as suitable benchmarks for evaluating the effectiveness of LLM unlearning. Some notable examples of these datasets include: The Enron dataset, which comprises employee emails publicly disclosed during Enron’s legal investigation by the Federal Energy Regulatory Commission (Wu et al., 2023b), the Training Data Extraction Challenge dataset used in (Jang et al., 2022), the Harry Potter book series dataset (Eldan & Russinovich, 2023; Shi et al., 2023), the toxicity generation dataset (Lu et al., 2022; Gehman et al., 2020), the TOFU dataset for unlearning fictitious entities (Maini et al., 2024), and the WMDP benchmark for accessing unlearning potential hazardous knowledge in domain of biology, cybersecurity, and chemistry (Li et al., 2024a). In what follows, we elaborate on the assessment of LLM unlearning in terms of unlearning effectiveness, utility preservation, and efficiency.
Unlearning effectiveness
@@
The efficacy of LLM unlearning can be examined from three perspectives: comparison with retraining (i.e., the gold standard of unlearning), ‘hard’ in-scope evaluation or robustness, and training data detection.
@@
LLM unlearning vs. retraining: In classic unlearning paradigms (Golatkar et al., 2020; Thudi et al., 2022a; Jia et al., 2023; Fan et al., 2024b), retraining a model from scratch after removing the forgotten data from the original training set is regarded as exact unlearning. However, the scalability challenges of retraining LLMs make it difficult to establish a performance upper bound for evaluating LLM unlearning. A recent solution in (Maini et al., 2024) is to incorporate fictitious data (synthetic author profiles) into the model training paradigm. Since the injected set never appeared in the original pretraining set, LLM fine-tuning can simulate the retraining process over the newly-introduced set. Another solution is to use a surrogate unseen forget set from a domain close to the domain of the real forget set to approximate a retrained model’s performance on the real forget data (Yao et al., 2024). Despite the progress with regard to the specialized TOFU dataset (Maini et al., 2024) and the method that approximates a retrained model’s performance, there is still a general need for precisely assessing the gap between (approximate) LLM unlearning methods and exact unlearning.
@@
‘Hard’ in-scope evaluation or robustness: As demonstrated in Sec. 3, unlearning is generally context and task-dependent, contingent upon an unlearning scope. Another effectiveness metric of LLM unlearning is to ensure forgetting concerning in-scope unlearned examples, even for those ‘hard’ ones that fall within the unlearning scope but may not be directly associated with the unlearning targets. The assessment of ‘hard’ in-scope examples can be achieved by techniques such as paraphrasing what LLMs intend to forget or creating multi-hop questions (Zhong et al., 2023). Evaluating ‘hard’ in-scope examples aligns seamlessly with the underlying principles of ‘worst-case’ or ‘adversarial’ evaluation methods for unlearning (Zhang et al., 2023d; Yong et al., 2023; Patil et al., 2024; Lynch et al., 2024; Fan et al., 2024a; Zhao et al., 2024). For instance, it is shown in (Yong et al., 2023) that unlearning a scope using an English-only example would not guarantee a similar unlearned outcome when translated into other languages. It is also crucial to evaluate the robustness of unlearned LLMs after fine-tuning. Recent studies have revealed that fine-tuning LLMs can sometimes lead to the re-emergence of behaviors that were not anticipated (Yang et al., 2023; Qi et al., 2023; Lermen et al., 2023; Yong et al., 2023).
@@
Training data detection, membership inference and data forging attacks: Membership inference attacks (MIA) (Shokri et al., 2017), designed to detect if a data point is part of a victim model’s training set, serve as a crucial data privacy-unveiled metric for evaluating machine unlearning methods (Thudi et al., 2022a; Jia et al., 2023). This metric gains even more significance in the context of LLM unlearning, particularly when retraining is not an option. This concept is also connected to training data memorization (Carlini et al., 2022), as well as training data extraction attacks (Nasr et al., 2023) in LLMs. However, evidence shows that existing state-of-the-art MIA methods for LLMs are limited in their ability to effectively distinguish membership and non-membership (Duan et al., 2024), suggesting opportunities for further research. In the realm of LLM unlearning, other privacy-related evaluation metrics have been explored and considered in various studies (Shi et al., 2023; Wu et al., 2023b; Jang et al., 2022; Pawelczyk et al., 2023; Maini et al., 2024; Zhang et al., 2024a).
Another important branch that affects the evaluation of efficacy of unlearning is data forging attacks (Thudi et al., 2022b). In these attacks, an adversary may be able to replace mini-batches used in training with different ones that yield nearly identical model parameters. These attacks may enable the claim of successful unlearning without actually unlearning samples while claiming they have been erased. These attacks are still under scrutiny and in (Suliman et al., 2024), Suliman et al. show that the errors associated with them may differ from model training with real (non-forged) data. The more developments in this area are needed to ensure verification of unlearning methods is effective and can be widely trusted.
Utility preservation
@@
Another crucial metric is to ensure the retained generation capabilities of unlearned LLMs on standard language modeling tasks that fall outside the unlearning scope. For example, evaluation on natural language understanding tasks (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023; Barbulescu & Triantafillou, 2024; Liu et al., 2024) and perplexity (Ilharco et al., 2022; Zhang et al., 2023c) has been considered in the literature. In line with evaluating the effectiveness of LLM unlearning on ‘hard’ in-scope examples, it is equally crucial to assess utility preservation using ‘hard’ out-of-scope examples, achieved by e.g., using data transformations or increasing the diversity of utility-oriented tasks. An example of ‘hard’ out-of-scope example is to use a retain set closely related to the domain of the unlearning target (Li et al., 2024a; Liu et al., 2024) to evaluate the unlearned model (e.g., unlearning economics while retaining econometrics). Lastly, we note that it can be difficult to determine the exact scope for some unlearning target (Hase et al., 2023b; Cohen et al., 2023), so part of the challenge here is deciding which generation capabilities should be retained in the first place.
Efficiency
@@
Computation cost has been a predominant efficiency metric when evaluating LLM unlearning methods, as demonstrated in Table 1. In addition to that, efforts have been made to extend LLM unlearning to black-box models, without access to model parameters, as demonstrated in (Pawelczyk et al., 2023). Furthermore, memory efficiency could also serve as a crucial efficiency metric. The distinction from parameter efficiency is that current parameter-efficient fine-tuning methods still impose substantial memory costs for storing LLMs and for executing back-propagation (Malladi et al., 2023). Thus, a future research direction is to explore memory-efficient fine-tuning methods for LLM unlearning.